OpenStudy (anonymous):

2 in a half dozen whole wheat-wheat muffins requires 600 g of flour. How many muffins can be made with 900 g of flour?

OpenStudy (anonymous):

I set it up like stoichiometry and ratios. 60 muffins (60 muffins in 2.5 dozen: i.e 24x2 =48+12= 60 muffins. 60 muffins/600 g flour... which also means it's 10 grams of flour per muffin. so you can just multiply the number of grams you have which is 900 and and divide it by how many grams per muffin which is 10. 900/10 is 90 muffins.

OpenStudy (mathmale):

Don't you mean "two and one half dozen" instead of "2 in a half dozen?" How many are in "two and one half dozen?"

OpenStudy (anonymous):

yeah

OpenStudy (mathmale):

Write a ratio involving the amount of flower and a ratio involving how many muffins, and then set these two ratios equal to each other. Two different quantities of flour are mentioned. What are these quantities? Write their ratio.

OpenStudy (mathmale):

Just re-read the problem statement, Danny. Two measures of flour are given. What's their ratio?

OpenStudy (anonymous):

um

OpenStudy (anonymous):

\[12\frac{ 1}{ 2 }:600\]

OpenStudy (mathmale):

The measures of flour are 600 and 900 g. What's their ratio? Be certain to use the simplest meaning of "ratio" here.

OpenStudy (mathmale):

Danny: All you have to do is to write this ratio:\[\frac{ 600~g }{ 900~g }=\frac{ 6 }{ 9 }=\frac{ 2 }{ 3 }.\] Now you'll need to set up another ratio: this time, write the ratio of 2 !/2 dozen, or 24, muffins to x muffins, where x represents the number of muffins that we can make with 900 g of flour. Write the ratio of 24 to x here, please.

OpenStudy (mathmale):

Once you've done that, equate the ratio 2/3 (from above) to your new ratio. Can you solve this equation for x?